converge or diverge by what test

n! 2n 00 E nal h3 ZEU u N=1
Answer #1

n! 22 Mo an 81 n 3 n n= (n+1)! 2nt 21 2 :: An Anti = (n+1)3 lim (n+1)! 2n+/ (n+1) 3 23 anti Now, L= lim n! an an noo n 700 L=